If demand (Qd)=100 -14p; and supply (Qs)= 50+6p; then at what price will there be a surplus of 60 units?


1
Expert's answer
2021-10-25T17:52:26-0400

Solution:

First, derive equilibrium price and quantity:

At equilibrium: Qd = Qs

100 – 14p = 50 + 6p

100 – 50 = 6p + 14p

50 = 20p

P = 2.5

Equilibrium price = 2.5

Substitute in the Qd function to derive equilibrium quantity:

Qd = 100 – 14p = 100 – (14 "\times" 2.5) = 100 – 35 = 65

Equilibrium quantity = 65

Consumer surplus = ½ "\times" 65 "\times" (P – 2.5)

60 = ½ "\times" 65 "\times" (P – 2.5)

P = 4.35